# CodeWiki: Evaluating AI’s Ability to Generate Holistic Documentation for Large-Scale Codebases

### v4

@@ 10            Recursively process $m^{\prime}$;
Our key innovation is adaptive scalability through dynamic delegation. When module complexity exceeds single-pass capacity, agents delegate sub-modules to specialized sub-agents. Delegation criteria: code complexity metrics (cyclomatic complexity, nesting depth), semantic diversity (functionally distinct subcomponents), and context window utilization.
The recursive process follows bottom-up processing (Algorithm 1). Upon delegation, agents provide sub-module specifications, the module tree updates, and newly created leaves process recursively. This handles modules of any size while maintaining quality.
Cross-Module Reference Management
Maintaining coherence across boundaries requires sophisticated reference management. When agents encounter external components during traversal, our intelligent resolution system creates cross-references rather than duplicating content. A global registry tracks documented components and locations, enabling quick identification of documented components.
This ensures conciseness while preserving completeness, creating interconnected documentation reflecting the actual codebase structure, enabling easy navigation between related components.
3.3 Hierarchical Assembly and Documentation Synthesis
4 Benchmarking System - CodeWikiBench
Repository-level documentation evaluation presents unique challenges due to complexity and hierarchical organization. Unlike code generation assessable through automated metrics, documentation quality requires nuanced evaluation of completeness, accuracy, and coherence across abstraction levels. Recent research highlights traditional metrics’ limitations for code tasks [11, 14], while LLM-as-a-Judge paradigms offer new possibilities [48, 8, 16]. We develop a comprehensive framework synthesizing repository-specific rubrics and employing agentic assessment.
4.1 Hierarchical Rubric Generation
Figure 2: Example hierarchical evaluation rubric for the RAGFlow repository. The rubric mirrors the project’s architectural structure with weighted requirements at multiple levels. Leaf nodes represent specific requirements assessed by Judge Agent, while parent scores are computed through weighted aggregation. The hierarchical organization ensures comprehensive coverage from high-level architectural components down to specific implementation details.
Following [44], we synthesize evaluation rubrics tailored for each repository, illustrated in Figure 2. We collect official documentation from selected open-source projects and hierarchically parse into structured JSON format, leveraging directory structure and markdown syntax. This captures multi-level technical documentation from architectural overviews to implementation details.
A Rubric Generator Agent processes each structure with tool access to detailed contents, generating hierarchical rubrics mirroring repository functionality and aligning with actual documentation standards. To enhance reliability, we employ multiple independent generations using different model families. Final rubrics emerge from synthesizing these perspectives, reducing single-model biases. Analysis shows generated rubrics demonstrate high consistency: 73.65% semantic reliability and 70.84% structural reliability (Appendix C).
4.2 Documentation Assessment
4.3 Hierarchical Score Aggregation
Leaf Assessment and Reliability
$S(\ell)=\bar{s}=\frac{1}{m}\sum_{i=1}^{m}s_{i}$
Reliability quantified via standard deviation:
$\sigma_{\ell}=\sqrt{\frac{1}{m-1}\sum_{i=1}^{m}(s_{i}-\bar{s})^{2}}$
Hierarchical Propagation
$S(n)=\frac{\sum_{i=1}^{k}w(c_{i})\cdot S(c_{i})}{\sum_{i=1}^{k}w(c_{i})},\quad\sigma_{n}=\frac{\sqrt{\sum_{i=1}^{k}w(c_{i})^{2}\sigma_{c_{i}}^{2}}}{\sum_{i=1}^{k}w(c_{i})}$
Final Assessment
Final repository score $S(R)\in[0,1]$ with deviation $\sigma_{R}$ represents comprehensive weighted aggregation propagating assessments and reliability upward. Lower $\sigma_{R}$ indicates higher methodology reliability. This dual-metric approach provides quality assessment with confidence bounds, enabling nuanced interpretation. Weights reflect component importance based on criticality and complexity while maintaining statistical rigor in uncertainty quantification.
Repository Language LOC System Score (%) Coverage Improvement (%)
All-Hands-AI–OpenHands Python 229,909 DeepWiki 73.04 ± 2.54 54/67
CodeWiki-kimi-k2 80.20 ± 3.26 60/67 +7.16
CodeWiki-sonnet-4 82.45 ± 2.65 59/67 +9.41
sveltejs–svelte JavaScript 124,576 DeepWiki 68.51 ± 3.31 76/96
CodeWiki-kimi-k2 70.41 ± 3.39 80/96 +1.90
CodeWiki-sonnet-4 71.96 ± 3.73 80/96 +3.45
puppeteer–puppeteer TypeScript 136,302 DeepWiki 64.46 ± 3.72 60/82
CodeWiki-kimi-k2 81.33 ± 3.33 67/82 +16.87
CodeWiki-sonnet-4 83.00 ± 3.37 74/82 +18.54
Unity-Technologies–ml-agents C# 86,106 DeepWiki 74.80 ± 3.69 39/46
CodeWiki-kimi-k2 69.64 ± 5.40 40/46 -5.16
CodeWiki-sonnet-4 79.78 ± 5.02 42/46 +4.98
elastic–logstash Java 117,485 DeepWiki 54.80 ± 4.10 38/57
CodeWiki-kimi-k2 56.89 ± 2.96 37/57 +2.09
CodeWiki-sonnet-4 57.90 ± 3.43 38/57 +3.10
wazuh–wazuh C 1,446,730 DeepWiki 68.68 ± 4.74 39/46
CodeWiki-kimi-k2 58.15 ± 6.20 32/46 -10.53
CodeWiki-sonnet-4 64.17 ± 5.44 34/46 -4.51
electron–electron C++ 184,234 DeepWiki 44.10 ± 3.12 54/92
CodeWiki-kimi-k2 36.99 ± 3.98 43/92 -7.11
CodeWiki-sonnet-4 42.30 ± 3.26 48/92 -1.80
Average DeepWiki 64.06 ± 3.60
CodeWiki-kimi-k2 64.80 ± 4.07 +0.74
CodeWiki-sonnet-4 68.79 ± 3.84 +4.73
Table 1: Documentation generation performance comparison across diverse repositories. Scores represent weighted averages with standard deviations indicating assessment consensus. Coverage shows satisfied leaf criteria out of total requirements. Blue-highlighted rows show CodeWiki-sonnet-4 results, while yellow-highlighted rows show CodeWiki-kimi-k2 results. Improvement column shows relative gain over DeepWiki baseline.
5 Experiment
We conduct comprehensive experiments to evaluate CodeWiki’s effectiveness in generating repository-level documentation across diverse programming languages and project types. Our evaluation is structured around three core research questions that address the fundamental challenges of automated repository documentation generation.
5.1 Research Questions
RQ2: Cross-Language Generalization Does CodeWiki demonstrate consistent performance across diverse programming languages, particularly comparing high-level scripting languages (Python, JavaScript, TypeScript) versus systems programming languages (C, C++, C#, Java)?
5.2 Baseline
CodeWiki provides an open-source framework specifically designed for comprehensive repository-level documentation generation. Due to the limited availability of comparable open-source solutions, we compare with closed-source alternatives. We select DeepWiki [10] as our primary baseline, as it shares our goal of automated repository-level documentation generation and has demonstrated effectiveness in industrial applications.
5.3 Experimental Setup
5.3.1 Repository Selection
We carefully curate a comprehensive benchmark consisting of 21 open-source repositories that represent diverse programming languages, project scales, and application domains. The selection criteria prioritize:
  1. 1.
    Language diversity to test cross-language generalization
  2. 2.
    Varying codebase sizes from 16K to 1.7M lines of code
  3. 3.
    Active maintenance with comprehensive official documentation for rubric generation
  4. 4.
    Diverse application domains including web frameworks, automation tools, machine learning platforms, and system utilities
Table 1 presents the main evaluation set of 7 repositories with detailed analysis, while Table 3 in Appendix D provides results for 14 additional repositories that further validate our approach across diverse codebases.
5.3.2 Model Configuration
We evaluate CodeWiki with two different LLM backends:
  • CodeWiki-sonnet-4: Uses Claude Sonnet 4 (proprietary model)
  • CodeWiki-kimi-k2: Uses Kimi K2 Instruct (open-source model)
For rubric generation, we employ multiple models (Claude Sonnet 4, Gemini 2.5 Pro, Kimi K2 Instruct) to reduce single-model bias. The evaluation process uses three independent judge models (Gemini 2.5 Flash, GPT OSS 120B, Kimi K2 Instruct) for robust assessment. Additional implementation details are provided in Appendix E.
5.3.3 Evaluation Protocol
For each repository, we execute the complete CodeWiki pipeline and generate comprehensive documentation including architectural overviews, component descriptions, and usage guides. The evaluation process follows our hierarchical rubric methodology:
  1. 1.
    Rubric Generation: Repository-specific rubrics derived from official documentation
  2. 2.
    Agentic Assessment: Multiple judge agents evaluate generated documentation against rubric criteria
  3. 3.
    Score Aggregation: Weighted scores computed with uncertainty quantification through standard deviation tracking
5.4 Results and Analysis
5.4.1 RQ1: Documentation Quality and Coverage
Table 1 presents comprehensive evaluation results addressing documentation quality. CodeWiki-sonnet-4 achieves an average quality score of 68.79%, representing a 4.73 percentage point improvement over DeepWiki’s 64.06%. Notably, CodeWiki-kimi-k2 demonstrates competitive performance at 64.80%, closely matching DeepWiki despite using an open-source model. This validates our framework’s effectiveness across different LLM backends.
The coverage metrics reveal that CodeWiki consistently satisfies more leaf-level requirements than the baseline. CodeWiki-sonnet-4 outperforms DeepWiki in 5 out of 7 repositories, with particularly strong improvements in TypeScript (83.00% vs. 64.46%, +18.54%), Python (82.45% vs. 73.04%, +9.41%), and JavaScript (71.96% vs. 68.51%, +3.45%). CodeWiki-kimi-k2 shows notable gains in TypeScript (+16.87%) and Python (+7.16%), demonstrating that our hierarchical approach enables strong performance even with open-source models.
Additional evaluation on 14 diverse repositories in Appendix D further validates these findings, with CodeWiki-kimi-k2 achieving an average improvement of 1.25% and substantial gains on large-scale projects including Trino (+29.53%) and Storybook (+11.42%).
Answer to RQ1: CodeWiki demonstrates strong overall documentation quality and coverage compared to existing systems. With proprietary models, it achieves 4.73% improvement; with open-source models, it matches baseline performance while offering transparency and accessibility benefits.
5.4.2 RQ2: Cross-Language Generalization
Our analysis reveals distinct performance patterns across programming language categories, as visualized in Figure 3.
Figure 3: Performance comparison by programming language categories. CodeWiki-sonnet-4 shows strong improvements in high-level scripting languages (+10.47%) and managed/enterprise languages (+4.04%), while systems programming languages present challenges for both systems.
CodeWiki-sonnet-4 demonstrates strong performance on high-level scripting languages, achieving 79.14% average across Python, JavaScript, and TypeScript repositories—a 10.47 percentage point improvement over DeepWiki’s 68.67%. CodeWiki-kimi-k2 achieves 77.31% on these languages, showing strong performance with open-source models (+8.64% over DeepWiki).
For managed/enterprise languages (C#, Java), CodeWiki-sonnet-4 achieves competitive performance with an average improvement of 4.04% (68.84% vs. 64.80%), while CodeWiki-kimi-k2 shows 63.27% (-1.53%). Systems programming languages present unique challenges where both frameworks exhibit similar patterns: CodeWiki-sonnet-4 achieves 53.24% compared to DeepWiki’s 56.39%, while CodeWiki-kimi-k2 obtains 47.58%. This suggests that specialized parsing strategies may benefit all approaches for handling complex constructs like pointer operations and template metaprogramming.
Answer to RQ2: CodeWiki demonstrates strong cross-language generalization with high performance on high-level languages (79.14% with proprietary models, 77.31% with open-source models) and consistent improvements on managed languages, while systems programming languages present similar challenges across all frameworks and model types.
5.4.3 RQ3: Scalability and Reliability
Repository size analysis demonstrates CodeWiki’s scalability across 86K-1.4M LOC. For high-level languages, CodeWiki maintains strong performance across scales: Unity ML-Agents (86K LOC) achieves 79.78%, while OpenHands (230K LOC) reaches 82.45%, demonstrating that hierarchical decomposition successfully addresses context limitations. Both CodeWiki and DeepWiki exhibit similar behavioral patterns across repository sizes within language families, with performance variations primarily attributable to language-specific parsing complexity rather than scale. Dynamic delegation automatically adapts to complexity levels, maintaining bounded processing requirements.
Reliability analysis through standard deviations (3.84% CodeWiki vs. 3.60% DeepWiki) validates our multi-model assessment methodology. Lower deviations for high-level languages (Python: 2.65%, JavaScript: 3.73%) versus systems languages (C: 5.44%, C++: 3.26%) confirm evaluation consistency correlates with language characteristics rather than repository scale.
Answer to RQ3: CodeWiki’s hierarchical decomposition enables consistent scalability across repository sizes from 86K to 1.4M LOC, with performance variations primarily attributable to language-specific characteristics rather than scale limitations. Both frameworks exhibit similar patterns across different repository sizes within language families.
6 Conclusion
We presented CodeWiki, an open-source semi-agentic framework addressing automated repository-level documentation through hierarchical decomposition and recursive processing. CodeWiki achieves 68.79% quality with proprietary models (Claude Sonnet 4) and 64.80% with open-source alternatives (Kimi K2 Instruct), both outperforming or matching DeepWiki’s 64.06%, demonstrating strong performance regardless of underlying model choice.
Results show performance variations correlate primarily with language characteristics rather than repository size or model type. High-level languages show consistent improvements across scales and models, while systems languages face challenges regardless of configuration.
Future directions include: developing specialized parsing modules for improved systems language support, extending to multi-version documentation tracking codebase evolution, and leveraging comprehensive documentation for downstream tasks like code migration. By releasing CodeWiki as open source with support for both proprietary and open-source LLMs, we accelerate research and enable the community to build upon our hierarchical processing approach with accessible technology choices.
Limitations
While CodeWiki demonstrates strong performance across multiple languages and repository scales, two main limitations warrant discussion. First, our framework exhibits reduced effectiveness on systems programming languages (C, C++) compared to high-level languages, achieving 53.24% versus 79.14% average scores. This gap stems from the inherent complexity of low-level constructs such as pointer operations, manual memory management, and template metaprogramming, which challenge both AST parsing and semantic understanding.
Second, our evaluation methodology, while more comprehensive than traditional metrics, relies on LLM-based assessment which may introduce model-specific biases despite our multi-model consensus approach. The average standard deviation of 3.84% indicates reasonable consistency, yet edge cases may benefit from human expert validation.
Appendices
Appendix A Example Generated Documentation
Appendix B Related Work
B.1 Code Intelligence and LLMs
The rapid advancement of Large Language Models (LLMs) has revolutionized code intelligence tasks. Early work by von der Mosel et al. [47] demonstrated the effectiveness of Transformer-based models for software engineering, leading to specialized code models like CodeBERT [12], GraphCodeBERT [13], CodeT5 [49], CodeT5+ [50], and CodeTF [3]. These established early transformer-based approaches for code understanding and generation.
Comprehensive surveys by Niu et al. [34] and Hou et al. [19] established the importance of domain-specific pre-training, showing consistent improvements over general-purpose language models across code search, summarization, and generation tasks. However, recent studies highlight important limitations: Li et al. [22] found that models often make confident decisions without sufficient reasoning, while Zheng et al. [57] noted that LLM performance remains unstable and requires careful prompt engineering [42].
For code generation, HumanEval [6] and MBPP [2] established foundational benchmarks, later extended to multiple languages by MultiPL-E [4] and CodeScope [52]. More recently, CodeMMLU [32] has emerged as a comprehensive multi-task benchmark for assessing code understanding and reasoning capabilities of CodeLLMs.
B.2 Automated Documentation Generation
Documentation generation has emerged as a critical LLM application in software engineering. Traditional approaches focused on function-level summarization using datasets like CodeSearchNet [15] and FunCom [21]. While Sun et al. [45] found that ChatGPT can generate coherent comments, their quality often falls short of specialized models when evaluated using BLEU and ROUGE metrics.
Recent research has shifted toward repository-level documentation generation, which requires understanding complex dependencies and architectural context. Luo et al. [27] introduced RepoAgent, an LLM-powered framework that leverages static analysis and incremental processing. DocAgent [53] proposed a multi-agent system with specialized roles (Reader, Searcher, Writer, Verifier) for collaborative documentation generation. Both systems demonstrate significant improvements over baseline approaches but struggle with scaling to large repositories and maintaining cross-file consistency.
Evaluation methodologies have evolved from traditional NLP metrics like BLEU [35] and ROUGE [23] to more sophisticated approaches. Roy et al. [41] and Mastropaolo et al. [29] questioned the effectiveness of automatic metrics, while recent work has adopted LLM-as-a-judge methodologies [56, 25]. Multi-dimensional evaluation frameworks now assess completeness, helpfulness, and truthfulness, recognizing that documentation quality cannot be reduced to a single dimension.
B.3 Multi-Agent Systems for Software Engineering
Multi-agent architectures have gained prominence for tackling complex software engineering tasks requiring diverse expertise and collaborative problem-solving. AutoGen [51] introduced a framework enabling multi-agent communication with language models, allowing agents to assume different roles and collaborate on complex tasks. MetaGPT [18] demonstrated meta-programming for collaborative frameworks, where agents follow software development methodologies to produce comprehensive solutions.
Multi-agent approaches include HyperAgent [36], which presents generalist software engineering agents capable of solving coding tasks at scale, and AgileCoder [33], which introduces dynamic collaborative agents for software development based on agile methodology. In the code generation domain, MapCoder [55] employed map-reduce style collaboration to break down complex programming problems, while ChatDev [38] showed that multi-agent systems can manage the entire software development lifecycle. The integration of multi-agent systems with documentation represents a recent innovation, with DocAgent demonstrating how role-based collaboration can improve documentation quality through specialized agents for analysis, information retrieval, drafting, and verification. However, existing approaches often lack systematic dependency management and struggle with maintaining consistency across large codebases.
Appendix C Rubric Reliability Assessment
@@
C.1 Semantic Consistency Analysis
@@
C.2 Structural Consistency Analysis
@@
C.3 Overall Reliability Score
@@
Repository Semantic (%) Structural (%)
All-Hands-AI–OpenHands 73.64 68.82
sveltejs–svelte 75.56 68.88
puppeteer–puppeteer 72.97 76.36
Unity-Technologies–ml-agents 75.95 71.22
elastic–logstash 73.13 68.22
wazuh–wazuh 72.20 67.62
electron–electron 72.12 74.79
Average 73.65 70.84
Appendix D Complementary Experimental Results
To further validate CodeWiki’s effectiveness across diverse programming languages and repository types, we conducted additional experiments on 14 repositories spanning 7 programming languages. Table 3 presents comprehensive results using CodeWiki-kimi-k2 (documentation generated using Kimi K2 Instruct) compared against the DeepWiki baseline.
Repository Lang. LOC System Score (%) Cov. Imp. (%)
chartjs–Chart.js JavaScript 90,925 DeepWiki 88.71 ± 3.29 57/61
CodeWiki-kimi-k2 84.01 ± 3.55 54/61 -4.70
marktext–marktext JavaScript 61,574 DeepWiki 81.64 ± 4.67 50/57
CodeWiki-kimi-k2 82.49 ± 2.51 51/57 +0.85
storybookjs–storybook TypeScript 364,606 DeepWiki 72.07 ± 3.59 64/78
CodeWiki-kimi-k2 83.49 ± 2.31 69/78 +11.42
mermaid-js–mermaid TypeScript 157,160 DeepWiki 73.81 ± 4.05 59/67
CodeWiki-kimi-k2 71.52 ± 5.36 57/67 -2.29
qmk–qmk_firmware C 1,099,824 DeepWiki 81.23 ± 2.62 94/110
CodeWiki-kimi-k2 72.01 ± 2.63 88/110 -9.22
sumatrapdfreader–sumatrapdf C 1,699,044 DeepWiki 55.81 ± 6.29 31/40
CodeWiki-kimi-k2 58.92 ± 5.18 31/40 +3.11
x64dbg–x64dbg C++ 207,671 DeepWiki 64.42 ± 2.07 62/107
CodeWiki-kimi-k2 58.01 ± 3.58 56/107 -6.41
nlohmann–json C++ 119,542 DeepWiki 66.06 ± 3.08 33/57
CodeWiki-kimi-k2 61.28 ± 2.35 30/57 -4.78
FluentValidation–FluentValidation C# 16,452 DeepWiki 76.48 ± 2.33 41/48
CodeWiki-kimi-k2 70.03 ± 2.90 38/48 -6.45
git-ecosystem–git-credential-manager C# 40,812 DeepWiki 87.03 ± 2.58 66/70
CodeWiki-kimi-k2 82.78 ± 0.90 61/70 -4.25
material-components–material-components-android Java 244,092 DeepWiki 70.99 ± 2.92 79/96
CodeWiki-kimi-k2 79.16 ± 4.32 89/96 +8.17
trinodb–trino Java 1,535,202 DeepWiki 55.27 ± 3.76 48/68
CodeWiki-kimi-k2 84.80 ± 2.84 62/68 +29.53
RasaHQ–rasa Python 147,711 DeepWiki 80.94 ± 3.15 64/70
CodeWiki-kimi-k2 81.10 ± 2.69 64/70 +0.16
microsoft–graphrag Python 30,390 DeepWiki 76.34 ± 3.28 48/56
CodeWiki-kimi-k2 79.29 ± 2.96 50/56 +2.95
Average DeepWiki 73.62 ± 3.41
CodeWiki-kimi-k2 74.87 ± 3.15 +1.25
Table 3: Complementary documentation generation performance evaluation across 14 additional repositories. Results demonstrate CodeWiki’s consistent effectiveness across diverse codebases, with notable improvements in Java repositories (trinodb–trino: +29.53%, material-components-android: +8.17%) and TypeScript (storybookjs–storybook: +11.42%).
These complementary results reinforce our findings from the main experiments, showing an average improvement of 1.25% across 14 diverse repositories. Notably, CodeWiki achieves substantial gains in several large-scale projects, including a 29.53% improvement on the Trino database engine (1.5M LOC) and 11.42% on the Storybook UI development platform (365K LOC). The results further validate that CodeWiki’s hierarchical approach generalizes well across different repository characteristics and programming paradigms.
Appendix E Implementation Details
E.1 Model Configurations

### v5

@@ 10            Recursively process $m^{\prime}$;
Appendix D Rubric Reliability Assessment
@@
D.1 Semantic Consistency Analysis
@@
D.2 Structural Consistency Analysis
@@
D.3 Overall Reliability Score
@@
Repository Semantic (%) Structural (%) All-Hands-AI–OpenHands 73.64 68.82 sveltejs–svelte 75.56 68.88 puppeteer–puppeteer 72.97 76.36 Unity-Technologies–ml-agents 75.95 71.22 elastic–logstash 73.13 68.22 wazuh–wazuh 72.20 67.62 electron–electron 72.12 74.79 Average 73.65 70.84
Appendix E Temporal Separation Details
Table 4 presents the specific commit information for each benchmark repository, demonstrating temporal separation from model training data.
Table 4: Repository commit information demonstrating temporal separation from model training data.
Repository Language Commit ID Commit Date OpenHands Python 30604c4 Sep 13, 2025 svelte JavaScript be645b4 Aug 31, 2025 puppeteer TypeScript c1105f1 Aug 29, 2025 ml-agents C# 4cf2f49 Aug 15, 2025 logstash Java 895cfa5 Aug 29, 2025 wazuh C 44b7cd3 Aug 30, 2025 electron C++ 828fd59 Aug 30, 2025
Appendix F Human Evaluation Pilot Study
To provide additional validation of our automated evaluation methodology beyond multi-model consensus, we conducted a pilot human study comparing CodeWiki and DeepWiki documentation outputs.
F.1 Study Design
Participants
We recruited three participants with diverse professional backgrounds relevant to documentation consumption:
  • Business Analyst: Represents stakeholders seeking high-level architectural understanding
  • AI Researcher: Represents technical users evaluating documentation for research purposes
  • Software Engineer: Represents developers using documentation for implementation guidance
None of the participants were prior contributors to the evaluated repositories, simulating the common scenario where developers approach unfamiliar codebases.
Evaluation Protocol
For each of three repositories (OpenHands, svelte, puppeteer), participants:
  1. 1.
    Received access to the official codebase and its official documentation
  2. 2.
    Reviewed anonymized documentation outputs from both CodeWiki and DeepWiki (presentation order randomized)
  3. 3.
    Assessed which documentation better reflected: (a) correctness relative to official documentation, and (b) coverage of actual repository architecture
  4. 4.
    Provided a binary preference judgment
The evaluation was conducted blind, with system identities hidden from participants.
F.2 Results
Table 5: Pilot human evaluation results showing participant preferences between CodeWiki (CW) and DeepWiki (DW) across three repositories.
Repository BA AI SE CW Pref.
OpenHands CW CW CW 3/3
svelte CW CW DW 2/3
puppeteer CW DW CW 2/3
Total 3/3 2/3 2/3 7/9 (77.8%)
Table 5 presents the evaluation results. CodeWiki was preferred in 7 out of 9 assessments (77.8%). Key observations:
Alignment with Automated Scores
Human preferences correlate with the magnitude of automated evaluation improvements:
  • OpenHands (+9.41% automated improvement): Unanimous CodeWiki preference (3/3)
  • puppeteer (+18.54% automated improvement): Strong CodeWiki preference (2/3)
  • svelte (+3.45% automated improvement): Closer preferences (2/3 for CodeWiki)
This alignment between human judgment and automated evaluation supports the validity of our LLM-based assessment methodology.
F.3 Limitations
We acknowledge several limitations of this pilot study:
  • Scale: Three participants and three repositories provide preliminary validation but not statistical significance
  • Simplified Protocol: Due to budget constraints, we used preference judgment rather than full rubric-based evaluation
  • Repository Selection: The evaluated repositories may not represent the full diversity of our benchmark
Comprehensive human evaluation with larger participant pools, more repositories, and detailed rubric-based assessment represents important future work to further validate the benchmark methodology.
Appendix G Implementation Details
G.1 Model Configurations